OpenStudy (cedric):

Simplify the following expression. 8x - 10y + 5z - 5x - 16y + 6z 3x - 6y + 11z 13x - 6y + 11z 13x - 26y + 11z 3x - 26y + 11z

OpenStudy (anonymous):

Do u know u can add like terms

OpenStudy (anonymous):

you need to find the slope here [if you are given the slope you know m as in y=mx+b] slope =rise/run [change in y]/[change in x] (y1-y2)/(x1-x2) either point can be 1 or 2 whichever is easier to subtract (8-2)/(-5+3) 6/-2 slope= -3 so you start with y= -3x to intercept a point subtract the given values from their relative x and y variable (-3, 2) and (-5, 8) (y-2)= -3(x+3) y-2 = -3x-9 y= -3x-7 or y= -3x (y-8)= -3(x+5) y-8= -3x-15 y= -3x-7

OpenStudy (anonymous):

I dont know what r u doing @abayomi12

OpenStudy (anonymous):

8x - 10y + 5z - 5x - 16y + 6z 8x-5x-10y-16y+5z+6z 3x-26y+11z

OpenStudy (anonymous):